how would I write an equation for the following linear relationships:
(x5,y20),(x10,y30)(x15,y40)(x20,y50)(x25,y50)(x30,y50)

Answer by richwmiller(17219)

There is something wrong because y doesn't change in the last three.
that can't be a linear equation since y does change in the first three.
I think the last three y's should be 50 60 70.
10/5=slope
(y-y1)=m(x-x1)
(y-20)=m(x-5)
y-20=2x-10
y=2x+10